The given data represent the total compensation for 10 andomly selected CEOs and their companys stock performance in 2009 AnCompensation (millions of dollars) Return (%) Stock 26.39 11.99 19.07 13.79 12.08 12.01 25.83 14.93 17.85 14.06 5.84 29.94 32Critical Values for Correlation Coefficient 1l 0.997 0.950 0.878 0.811 0.754 0.707 0.666 0.632 0.602 4 10
Answer #1

Ans) Here we need to predict stock return based on CEO.So here stock return is dependent variable and CEO is independent variable. Let the regression line be : y_i=\beta_0+\beta_1*x_i + e_i , where i = 1(1)10 and errors e,S are normal variates with homoscedastic variance. Now here it is given the data and correlation coefficient r between the two variables. The relation between β1( the estimator of 61) and r is given by: \hat\beta_1=r*\frac{s_y}{s_x} , where s_y is the sample standard deviation of y and s_x is the sample standard deviation of x. Here Sy 25.1 4683 and = 5.458811 and -0.1876. So the value of 31 =一0.8642076 . Now \hat\beta_0 (the estimator of \beta_0 ) is given by : \hat\beta_0=\bar{y}-\beta_1*\bar{x} where \bar{y} is the sample mean of y and \bar{x} is the sample mean of x. Here y 17.228 and T16.8 .So the value of 30 31.7466871 .

Now the predicted line is given by: у + β1 * 2-31.7466871-0.8642076 * 2.

Now put 15 and we get У 18.78357 18.8 upto one decimal approximation  and put r25 and we get  y= 10.1415